# ProactiveBench: A Comprehensive Benchmark Evaluating Proactive Interactions in Video Large Language Models

### v1

@@ Abstract
With the growing research focus on multimodal dialogue systems, the capability for proactive interaction is gradually gaining recognition. As an alternative to conventional turn-by-turn dialogue, users increasingly expect multimodal systems to be more initiative, for example, by autonomously determining the timing of multi-turn responses in real time during video playback. To facilitate progress in this emerging area, we introduce ProactiveBench, the first comprehensive benchmark to evaluate a system’s ability to engage in proactive interaction. Since model responses are generated at varying timestamps, we further propose PAUC, the first metric that accounts for the temporal dynamics of model responses. This enables a more accurate evaluation of systems operating in proactive settings. Through extensive benchmarking of various baseline systems on ProactiveBench and a user study of human preferences, we show that PAUC is in better agreement with human preferences than traditional evaluation metrics, which typically only consider the textual content of responses. These findings demonstrate that PAUC provides a more faithful assessment of user experience in proactive interaction scenarios.  [^1] Project homepage: https://github.com/yellow-binary-tree/ProactiveBench
ProactiveBench: A Comprehensive Benchmark Evaluating Proactive Interactions in Video Large Language Models
@@ 1 Introduction
To address this gap and advance research on proactive interaction, we introduce ProactiveBench, the first comprehensive benchmark specifically designed to evaluate the proactive interaction capabilities of MLLMs. ProactiveBench specifically targets the task of question answering: given a question presented at the very beginning of the video, the system must proactively detect when relevant information appears in one or more video segments and initiate responses accordingly as the video progresses. Compared to more open-ended interactive tasks, this question-answering setup offers relatively objective evaluation criteria (e.g., the availability of ground-truth answers), making it particularly suitable for academic research. By constructing a suite of diverse tasks and collecting videos from a wide range of sources, ProactiveBench encompasses several representative application scenarios for proactive interaction. It includes a broad spectrum of video topics, integrates multiple modalities, and supports multi-turn outputs to reflect realistic and varied use cases.
@@
(1) Benchmark with Multi-turn, fully open-ended answers. Unlike most existing video question answering benchmarks that predominantly adopt multiple-choice formats, ProactiveBench features fully open-ended questions requiring multi-turn, free-form textual responses. While this design introduces additional challenges for evaluation, it offers a more realistic and comprehensive assessment of a model’s interactive capabilities and is closer to real-world application scenarios.
(2) Benchmark with diverse topics and multimodal inputs. ProactiveBench includes videos covering a wide array of topics that are highly relevant to proactive interaction, such as web videos, egocentric recordings, television series, and surveillance footage. It also incorporates multiple input modalities including text, video, and speech, reflecting the complexity and richness of real-world use cases.
@@ 3.1 Adjusting the Importance of Timeliness
4 The ProactiveBench Benchmark
To make use of PAUC, a benchmark is required for evaluating proactive models. We introduce ProactiveBench, the first comprehensive benchmark designed for proactive interaction. To encompass prevalent scenarios in proactive interaction, ProactiveBench focuses on four key tasks:
@@ 4.1.1 Data Source
There are already many offline VideoQA benchmarks focusing on various capabilities. To construct ProactiveBench, we source video and annotations from Shot2story-MAGQA-39k for [WEB] Wang et al. (2024); Han et al. (2023), Ego4D Goalstep Song et al. (2023) for [EGO], TVQA Lei et al. (2018) for [TV], and UCF-Crime Sultani et al. (2018); Yuan et al. (2023) for [VAD].
4.1.2 Question and Answers in ProactiveBench
@@
Table 1: Dataset Statistics for each task from ProactiveBench.
@@ 5 Employing Offline Video-Text LLMs for Proactive Interaction
To the best of our knowledge, as of the writing of this manuscript, only two video-text LLMs designed for proactive interaction, VideoLLM-Online Chen et al. (2024a) and MMDuet Wang et al. (2024), have fully open-sourced their code to support proactive evaluation. To enable the evaluation of a broader set of models on ProactiveBench, we employ a simple rule-based strategy to adapt offline video-text LLMs for proactive interaction. Specifically, we segment each video into fixed-length chunks and, at each timestep, provide the model with the current video chunk, the associated question, and the model’s previous response as input. The model is first required to determine whether the current video chunk can answer the question, and output one of the following: “I have no answer,” “I have the same answer” (as the previous response), or “I have a new answer.” If the model responds with “I have a new answer,” it is then required to generate the updated response accordingly. In practice, we find that only proprietary models are capable of reliably following these multi-step instructions. Open-source models, by contrast, typically fail to comply and result in unexpected behaviors (e.g., ignore the instructions and start answering questions directly). Consequently, for open-source models, we adopt a simplified strategy: each video chunk is presented to the model alongside the question, and the model is asked to determine whether the chunk contains sufficient information to answer the question. If the model responds affirmatively, we perform an additional round of inference to obtain the answer based on the current chunk.
@@ 6 Experiments
Table 3: Results on ProactiveBench with different $\omega$.
@@
We report PAUC metric on ProactiveBench for the following methods: (1) proprietary offline video MLLMs, (2) open-sourced offline video MLLMs, (3) open-sourced proactive video MLLMs, and (4) human performance.
@@ 7 Conclusion
In this work, we address the emerging challenge of evaluating proactive multimodal dialogue systems by introducing ProactiveBench, the first comprehensive benchmark for assessing systems’ capabilities in proactive interaction scenarios. We propose PAUC, a novel evaluation metric that tracks how the quality of a model’s responses evolves throughout the video, as a more valid metric to evaluate models for proactive interaction. We believe that ProactiveBench and PAUC will serve as valuable tools for future research and development in proactive interaction models that can have potentially high-impact application areas such as live stream understanding, video anomaly detection, and ego-centric agents.
@@ Appendix B Case Study
Figure 6: Qualitative demonstration of ProactiveBench and PAUC. The first example is from [WEB], and the second example is from [VAD]. The underlined content in the model responses highlights key information (for visualization purposes only, not as part of the evaluation criteria). In the line chart below, each dot represents a model response and its impact on the score poly-line.

### v2

# ProactiveVideoQA: A Comprehensive Benchmark Evaluating Proactive Interactions in Video Large Language Models

@@ Abstract
With the growing research focus on multimodal dialogue systems, the capability for proactive interaction is gradually gaining recognition. As an alternative to conventional turn-by-turn dialogue, users increasingly expect multimodal systems to be more initiative, for example, by autonomously determining the timing of multi-turn responses in real time during video playback. To facilitate progress in this emerging area, we introduce ProactiveVideoQA, the first comprehensive benchmark to evaluate a system’s ability to engage in proactive interaction. Since model responses are generated at varying timestamps, we further propose PAUC, the first metric that accounts for the temporal dynamics of model responses. This enables a more accurate evaluation of systems operating in proactive settings. Through extensive benchmarking of various baseline systems on ProactiveVideoQA and a user study of human preferences, we show that PAUC is in better agreement with human preferences than traditional evaluation metrics, which typically only consider the textual content of responses. These findings demonstrate that PAUC provides a more faithful assessment of user experience in proactive interaction scenarios.  [^1] Project homepage: https://github.com/yellow-binary-tree/ProactiveVideoQA
ProactiveVideoQA: A Comprehensive Benchmark Evaluating Proactive Interactions in Video Large Language Models
@@ 1 Introduction
To address this gap and advance research on proactive interaction, we introduce ProactiveVideoQA, the first comprehensive benchmark specifically designed to evaluate the proactive interaction capabilities of MLLMs. ProactiveVideoQA specifically targets the task of question answering: given a question presented at the very beginning of the video, the system must proactively detect when relevant information appears in one or more video segments and initiate responses accordingly as the video progresses. Compared to more open-ended interactive tasks, this question-answering setup offers relatively objective evaluation criteria (e.g., the availability of ground-truth answers), making it particularly suitable for academic research. By constructing a suite of diverse tasks and collecting videos from a wide range of sources, ProactiveVideoQA encompasses several representative application scenarios for proactive interaction. It includes a broad spectrum of video topics, integrates multiple modalities, and supports multi-turn outputs to reflect realistic and varied use cases.
@@
(1) Benchmark with Multi-turn, fully open-ended answers. Unlike most existing video question answering benchmarks that predominantly adopt multiple-choice formats, ProactiveVideoQA features fully open-ended questions requiring multi-turn, free-form textual responses. While this design introduces additional challenges for evaluation, it offers a more realistic and comprehensive assessment of a model’s interactive capabilities and is closer to real-world application scenarios.
(2) Benchmark with diverse topics and multimodal inputs. ProactiveVideoQA includes videos covering a wide array of topics that are highly relevant to proactive interaction, such as web videos, egocentric recordings, television series, and surveillance footage. It also incorporates multiple input modalities including text, video, and speech, reflecting the complexity and richness of real-world use cases.
@@ 3.1 Adjusting the Importance of Timeliness
4 The ProactiveVideoQA Benchmark
To make use of PAUC, a benchmark is required for evaluating proactive models. We introduce ProactiveVideoQA, the first comprehensive benchmark designed for proactive interaction. To encompass prevalent scenarios in proactive interaction, ProactiveVideoQA focuses on four key tasks:
@@ 4.1.1 Data Source
There are already many offline VideoQA benchmarks focusing on various capabilities. To construct ProactiveVideoQA, we source video and annotations from Shot2story-MAGQA-39k for [WEB] Wang et al. (2024); Han et al. (2023), Ego4D Goalstep Song et al. (2023) for [EGO], TVQA Lei et al. (2018) for [TV], and UCF-Crime Sultani et al. (2018); Yuan et al. (2023) for [VAD].
4.1.2 Question and Answers in ProactiveVideoQA
@@
Table 1: Dataset Statistics for each task from ProactiveVideoQA.
@@ 5 Employing Offline Video-Text LLMs for Proactive Interaction
To the best of our knowledge, as of the writing of this manuscript, only two video-text LLMs designed for proactive interaction, VideoLLM-Online Chen et al. (2024a) and MMDuet Wang et al. (2024), have fully open-sourced their code to support proactive evaluation. To enable the evaluation of a broader set of models on ProactiveVideoQA, we employ a simple rule-based strategy to adapt offline video-text LLMs for proactive interaction. Specifically, we segment each video into fixed-length chunks and, at each timestep, provide the model with the current video chunk, the associated question, and the model’s previous response as input. The model is first required to determine whether the current video chunk can answer the question, and output one of the following: “I have no answer,” “I have the same answer” (as the previous response), or “I have a new answer.” If the model responds with “I have a new answer,” it is then required to generate the updated response accordingly. In practice, we find that only proprietary models are capable of reliably following these multi-step instructions. Open-source models, by contrast, typically fail to comply and result in unexpected behaviors (e.g., ignore the instructions and start answering questions directly). Consequently, for open-source models, we adopt a simplified strategy: each video chunk is presented to the model alongside the question, and the model is asked to determine whether the chunk contains sufficient information to answer the question. If the model responds affirmatively, we perform an additional round of inference to obtain the answer based on the current chunk.
@@ 6 Experiments
Table 3: Results on ProactiveVideoQA with different $\omega$.
@@
We report PAUC metric on ProactiveVideoQA for the following methods: (1) proprietary offline video MLLMs, (2) open-sourced offline video MLLMs, (3) open-sourced proactive video MLLMs, and (4) human performance.
@@ 7 Conclusion
In this work, we address the emerging challenge of evaluating proactive multimodal dialogue systems by introducing ProactiveVideoQA, the first comprehensive benchmark for assessing systems’ capabilities in proactive interaction scenarios. We propose PAUC, a novel evaluation metric that tracks how the quality of a model’s responses evolves throughout the video, as a more valid metric to evaluate models for proactive interaction. We believe that ProactiveVideoQA and PAUC will serve as valuable tools for future research and development in proactive interaction models that can have potentially high-impact application areas such as live stream understanding, video anomaly detection, and ego-centric agents.
@@ Appendix B Case Study
Figure 6: Qualitative demonstration of ProactiveVideoQA and PAUC. The first example is from [WEB], and the second example is from [VAD]. The underlined content in the model responses highlights key information (for visualization purposes only, not as part of the evaluation criteria). In the line chart below, each dot represents a model response and its impact on the score poly-line.